OpenStudy (anonymous):

Express the function as the sum of a power series by first using partial fractions. f(x)= 13/(x^2 − 3x − 40) sum n=0 to infinity

OpenStudy (anonymous):

f(8) is undefined..

OpenStudy (anonymous):

?

OpenStudy (anonymous):

fortunately this problem has been cooked so the partial fractions are easy to find

OpenStudy (anonymous):

\[\frac{13}{(x-8)(x+3)}=\frac{A}{x-8}+\frac{B}{x+3}\] you know how to find the partial fractions?

OpenStudy (anonymous):

stupid typo \[\frac{13}{(x-8)(x+5)}=\frac{A}{x-8}+\frac{B}{x+5}\]

OpenStudy (anonymous):

you can eyeball them i can show you a snap way to do it if you like

OpenStudy (anonymous):

or did you find them already?

OpenStudy (anonymous):

sum n=0 to inifnity [(-5)^n+1 -(1/(8^n+1))]x^n is that correct??:S

OpenStudy (anonymous):

i kinda doubt it, maybe

OpenStudy (anonymous):

oh maybe it is right lets see

OpenStudy (anonymous):

ok.s

OpenStudy (anonymous):

\[\frac{1}{x-8}=-\frac{1}{8-x}=-\frac{1}{8}\frac{1}{1-\frac{x}{8}}\] \[=-\frac{1}{8}(1+\frac{x}{8}+\left(\frac{x}{8}\right)^2+...\]

OpenStudy (anonymous):

\[=-\sum(\frac{x}{8})^n\] if my algebra is correct

OpenStudy (anonymous):

the next one alternates \[\sum(-1)^n(\frac{x}{5})^n\]

OpenStudy (anonymous):

so your answer looks right except i think you made a typo

OpenStudy (anonymous):

typo ..where?